OpenStudy (anonymous):

Find the values of x for which f(x) = [x(x − 2)]2 is an increasing function. Also, find the points on the curve, where the tangent is parallel to x-axis.

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad \left[x(x-2)\right]^2\]So for the first part: In order to find intervals of increasing/decreasing, we need to find critical points. Do you remember how to do that? :o

zepdrix (zepdrix):

I assume this is for Calculus, yes? :)

OpenStudy (anonymous):

could you help me

zepdrix (zepdrix):

Can you answer my questions..?

OpenStudy (anonymous):

linear function

OpenStudy (amoodarya):

f'=2(x)(x-2)^2+2(x-2)x^2=2x(x-2)(x-2+x) f'=4x (x-1)(x-2) =0 x=0 , 1,2